OpenStudy (anonymous):

Use the Quotient Rule to differentiate the function f(x)=3x/(x^(4)+8)

OpenStudy (anonymous):

\[f(x)=\frac{3x}{x^4+8}=\frac{u}{v}\]\[f'(x)=\frac{v \times \frac{du}{dx} - u \times \frac{dv}{dx}}{v^2}\] \[u=3x \Rightarrow \frac{du}{dx} = ?\]\[v=x^4+8 \Rightarrow \frac{dv}{dx}=?\]